A geometric series that has 7 terms it starts with 1000 and decreases successively by 20 %

Answer by Edwin McCravy(20056)

To decrease by 20% is to be multiplied by 100%-20% or 80% or
multiplied by 0.80:

1000×0.80 = 800
800×0.80 = 640
640×0.80 = 512
512×0.80 = 409.6
409.6×0.80 = 327.68
327.68×0.80 = 262.144
262.144×0.80 = 209.7152

Edwin